OpenStudy (anonymous):

Help with the following integral: ∫(from 0 to pi/2) sin^2(x) cos^2(x) sorry I don't really know how to write integrals down on the computer :/

OpenStudy (anonymous):

I got to this point : sin^2(x) (1+cos^(x) /2 )

OpenStudy (experimentx):

change that into sin2x

OpenStudy (anonymous):

Well I can get 1/2 sin2x with (sinxcosx)^2 would that work?

OpenStudy (experimentx):

no ... there is square again.

OpenStudy (anonymous):

but (sinxcosx)^2 = (1/2sin2x)^2 = 1/4(sin^2(2x))?

OpenStudy (experimentx):

yes ... but still that will not work. there is square ... but we can change it into cos4x

OpenStudy (anonymous):

Yeah I see it.. using basic trig! thanks :)